OpenStudy (anonymous):

Simplify -3/4 (16x + 12y -8) A. –4x – 3y + 2 B. –12x – 9y – 8 C. –12x + 9y – 6 D. –12x – 9y + 6

OpenStudy (anonymous):

Is D the correct answer?

OpenStudy (anonymous):

yes the correct answer is D.

OpenStudy (anonymous):

do you want me to explain??

OpenStudy (anonymous):

\[-3/4 (16x+12y-8) = -3 (4x+3y-2) = -12x +-9y + 6\]

OpenStudy (anonymous):

No no I understand I just wanted to know if I did it correctly!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

THANKS A LOT!